Let QnR be a map defined by f())=f(V2) . Show that \phi is a ring homomorphism, and Im() is a field.

QnR
f())=f(V2)

Im()
Answer #1

: Ql R defined b P(f)f) Let fe, fdx)e Q [x] (fe+= ( f18)= f+8)(E) =f(TEJ+ &«IEJ = 4(ftesj+ ¢ (dr), (f)) CE P- = (a) (P) Lfe).中(fa) = 0} Long ={fe) e a x7: kenp 2 Now balroowjal fnae 22-2 is ecluci ble OVeN fietd Hence fr(ps a fieldd